# Asymmetric Errors

### v1

@@ Abstract
We present a procedure for handling asymmetric errors. Many results in particle physics are presented as values with different positive and negative errors, and there is no consistent procedure for handling them. We consider the difference between errors quoted using pdfs and using likelihoods, and the difference between the rms spread of a measurement and the 68% central confidence region. We provide a full analysis of the possibilities, and software tools to enable their use.
@@ 1.1 Background and motivation
Results in particle physics are often given with errors which are asymmetric, of the form ${x}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$. For instance, taking some results from the latest EPS conference:
  • ATLAS quote their measurement of the Higgs width as $\Gamma_{H}={4.5}^{+3.3}_{-2.5}$ MeV [ATLAShiggs].
  • CMS quote the same quantity as $\Gamma_{H}={3.2}^{+2.4}_{-1.7}$ MeV [CMShiggs].
  • NOvA have measured the neutrino CP violating parameter as $\delta_{CP}={0.82}^{+0.27}_{-0.87}\,\pi$ [NOvA].
  • Belle II quote the branching ratio for $B\to\rho^{+}\rho^{0}$ as $({23.2}^{+2.2}_{-2.1}(stat.)\pm 2.7(sys.))\times 10^{-6}$ [Belle].
  • LHCb give the difference in the decay width for the neutral $B$ mesons as $\Gamma_{s}-\Gamma_{d}=\linebreak{-0.0056}^{+0.0013}_{-0.0015}(stat.)\pm 0.0014(sys.)\,{\rm ps}^{-1}$ [LHCb].
Despite their widespread use, their interpretation and their handling is unclear and discussion in the literature is limited [Schmelling, Systematic, Statistical, dagostini, Possolo]. This paper explores the reasons why they appear, and gives methods for their consistent handling.
@@
    when systematic errors are being studied using the OPAT (“One Parameter At a Time”) method. Often a nuisance parameter is changed by ±$\pm\sigma$ to observe the shift this produces in a result, and, as illustrated in Figure 1, the upward and downward shifts are different. Alternatively, when it is technically possible, many values can be generated according to a Gaussian distribution, to find and quantify the distribution of the results, and this may not be symmetric.
@@
    from maximum likelihood estimation with errors given by the $\Delta\ln L=-{1\over 2}$ method, and the log likelihood is not a symmetric parabola. This is also shown in Figure 1.
@@
    when a random variable is not distributed according to a Gaussian distribution, such as a Poisson distribution with small $N$. This includes as an important special case the replacement of a Gaussian random variable by another which is a non-trivial transformation, as happens on a plot with a logarithmic scale.
@@
Figure 1: The two classes of asymmetric error, from OPAT systematic studies (left) and maximum likelihood estimation (right)
@@ 1.2 Beyond the Gaussian
for a measurement $x$ of some value $a$. This is what is encoded in any reported result of the form $a=x\pm\sigma$. When a result is reported as $a={x}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ then this is making a clear analogy with the Gaussian form. But this similarity conceals two important questions which cannot be evaded.
@@
There is no universal answer; this has to be done case by case. If one is presenting a final result, then the 68% confidence level (CL) definition is clearly preferable. One wants to make a statement about the true value of the parameter, not about one’s apparatus. On the other hand, for an intermediate quantity which is going to be used as part of the final result, one is going to want to combine errors in quadrature, and here one is likely to be handling rms errors: variances add, even for non-Gaussian distributions.
@@
We assert that the answers to these two questions are linked. If we are considering a pdf $p(x;a)$ then we can integrate it to obtain the spread (as given by the variance) of the measured $x$ for a given value of $a$ and the ‘errors’ are expression of that spread. It tells us nothing about $a$. If we are dealing with a likelihood function $L(a;x)$ we can get the ‘best’ estimate $\hat{a}$ and the 68% CL central region from the $\Delta\ln L=-{1\over 2}$ points, but we can say nothing about $x$. It is true that a pdf can give 68% central limits, but they are the horizontal lines on the Neyman confidence belt, not the vertical ones. So an ‘asymmetric error’ may refer to the variance and skewness of a pdf, or to the confidence region for a likelihood. Both cases arise in practice. We will therefor consider asymmetric pdfs and the variances (and higher moments) of $x$, and also, separately, asymmetric likelihoods and the confidence regions for $a$.
@@
     Usage varies in the literature, but within this paper we use the variable names $x$ and sometimes $y$ to refer to results of measurements, and $a$ to refer to an ideal ‘true’ value. Thus pdfs are considered as functions of $x$ and likelihoods as functions of $a$. Alternative names may be used in particular examples, as when $R$ and $\nu$ (for ’Result’ and ’nuisance’) are used in the OPAT plot of Figure 1 rather than $\hat{a}$ and $x$.
@@ 1.2.1 Combination of Errors
This simple formula is where most physicists first meet statistics, in the context of experimental errors and uncertainties. For some function $u=u(x,y)$ where $x$ and $y$ are independent and the errors are small (so that  linear approximation is valid),
@@
We need to extend it to cope with $\sigma^{+}_{x},\sigma^{-}_{x},\sigma^{+}_{y},\sigma^{-}_{y}$ and thus $\sigma^{+}_{f},\sigma^{-}_{f}$.
In introductory texts this formula is used in problems like determining the speed from the distance travelled in a given time using $v=x/t$, or the acceleration due to gravity from the length and period of a simple pendulum using $g=4\pi^{2}L/T^{2}$. In more advanced experimental work it is used in instances like calculating a branching ratio, $Br={N-b\over\eta}$, from some number $N$ of observed events, a background $b$ and an efficiency $\eta$, or in fitting a straight line (or some more sophisticated function) to a set of values $\{x_{i},y_{i}\pm\sigma_{i}\}$.
@@
where the two random variables $X_{1}$ and $X_{2}$ are supposed independent. If the pdfs are Gaussian, as shown in the left hand plot of Figure 2, where the dotted lines are the lines of constant $x_{2}-x_{1}$, then the integral gives a Gaussian for $\ell$ with standard deviation $\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}$, in accordance with Equation (2).
Figure 2: Combination of errors using pdfs (left) and likelihoods (right)
@@ 1.2.2 Combination of Results
But it is also possible that one has a set of results presented as pdfs: the $\hat{a}_{i}$ are estimators of the true $a$, they are functions of the data $x_{i}$ and are therefore described by pdfs. If they are unbiassed then any weighted sum $\sum_{i}w_{i}a_{i}$ is an unbiased estimate, and one has the freedom to choose the $w_{i}$ to minimise the variance on this, using the combination of errors formula, Equation (2). For Gaussian errors this gives the well known prescription $w_{i}={1/\sigma_{i}^{2}\over\sum_{j}1/\sigma_{j}^{2}}$: the weight is proportional to the inverse square of the error. Hence for Gaussian errors the combination results using pdfs and using likelihoods gives the same answer, even though the meaning of the answer is subtly different. Again, for non-Gaussian errors this is not the case and we must consider both separately.
1.3 Why the ‘usual procedure’ is wrong
The flaw in the logic is as follows: when two random variables, described by dimidiated Gaussians, are combined, there is a 25% chance that they will both fluctuate upwards. That distribution is indeed described (within the framework of this model) by a Gaussian whose standard deviation is the sum in quadrature of the two components. Likewise there is a 25% chance that they will both go negative, described by $\sqrt{{\sigma^{-}}_{1}^{2}+{\sigma^{-}}_{2}^{2}}$. But there is a 50% chance that one will go up and the other will go down. This is neglected by the ‘usual procedure’, and it is this filling in of central values that lets the Central Limit Theorem do its work.
Examples of how this can give wrong results are in Section 4.1.2
@@ 2.1 Modelling pdf errors
    the quantiles. The median $M$ is the 50% quantile $x_{0.50}$, and the 68% (one sigma) central confidence probability region $[x_{0.16},x_{0.84}]$ can be written as $[M-{\sigma^{-}},M+{\sigma^{+}}]$. This can be expressed in the form ${M}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ and will be referred to as “the quantile parameters” in what follows.
@@
It might be thought that the mean could be used as an alternative to the median in combination with the bounds of the 68% central confidence region, writing it as $[x_{0.16},x_{0.84}]=[\mu-{\sigma^{-}},\mu+{\sigma^{+}}]$ However this parameterisation has an unhelpful behaviour in some models, as will be noted in Sections A.8 and A.9, and it will not be pursued further.
@@
Figure 3: The dimidiated and distorted Gaussians for ${5.00}^{+1.0}_{-0.9}$ and ${5.00}^{+0.85}_{-1.15}$

### v2

@@ Abstract
We present a procedure for handling asymmetric errors. Many results in particle physics are presented as values with different positive and negative errors, and there is no consistent procedure for handling them. We consider the difference between errors quoted, using pdfs and using likelihoods, and the difference between the rms spread of a measurement and the 68% central confidence region. We provide a comprehensive analysis of the possibilities, and software tools to enable their use.
@@ 1.1 Background and motivation
Results in particle physics are often given with errors which are asymmetric, of the form ${x}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$. For instance, taking some results from the latest (at the time of writing) EPS conference:
  • ATLAS quote their measurement of the Higgs width as $\Gamma_{H}={4.5}^{+3.3}_{-2.5}$ MeV [1].
  • CMS quote the same quantity as $\Gamma_{H}={3.2}^{+2.4}_{-1.7}$ MeV [2].
  • NOvA have measured the neutrino CP violating parameter as $\delta_{CP}={0.82}^{+0.27}_{-0.87}\,\pi$ [3].
  • Belle II quote the branching ratio for $B\to\rho^{+}\rho^{0}$ as $({23.2}^{+2.2}_{-2.1}(stat.)\pm 2.7(sys.))\times 10^{-6}$ [4].
  • LHCb give the difference in the decay width for the neutral $B$ mesons as $\Gamma_{s}-\Gamma_{d}=\linebreak{-0.0056}^{+0.0013}_{-0.0015}(stat.)\pm 0.0014(sys.)\,{\rm ps}^{-1}$ [5].
Despite their widespread use, their interpretation and their handling is unclear and discussion in the literature is limited [6, 7, 8, 9, 10, 11].This paper explores the reasons why they appear, and gives methods for their consistent handling. It addresses the basic question of how to use results presented as ${R}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$, in the absence of any further information.
@@
    when systematic errors are being studied, often using the OPAT (“One Parameter At a Time”) method when profiling out the nuisance parameters in the full likelihood is not feasible. In evaluating a systematic error due to the uncertainty (which can be Bayesian or frequentist) of a nuisance parameter $\nu$ one can change itby ±$\pm\sigma$ to observe the shift this produces in a result, and, as illustrated in Figure 1, the upward and downward shifts are different. Alternatively, when it is technically possible, many values can be generated according to a Gaussian distribution, to find and quantify the distribution of the results, and this may not be symmetric.
@@
    from maximum likelihood estimation with errors given by the $\Delta\ln L=-{1\over 2}$ method, and the log likelihood is not a symmetric parabola. This is also shown in Figure 1. (The $\Delta\ln L=-{1\over 2}$ method is an approximation, but its use is widespread, essentially universal, and in what follows this is what is meant by ’likelihood errors’.)
@@
    when a random variable is distributed according to a non-Gaussian distribution, such as a Poisson distribution with small mean. This includes as an important special case the replacement of a Gaussian random variable by another variable according to a non-trivial transformation, as happens on a plot with a logarithmic scale.
@@
Figure 1: The two classes of asymmetric error, from OPAT systematic studies of changes to a result $R$ in response to a nuisance parameter $\nu$ (left) and maximum likelihood estimation for a parameter $a$ (right).
     For the avoidance of confusion we point out that in this paper ${\sigma^{-}}$ is positive, like ${\sigma^{+}}$. And if $R$ falls with $\nu$ (rather than rising as shown in Figure 1), then ${\sigma^{+}}$ and ${\sigma^{-}}$ are still the magnitudes of the positive and negative deviations in $R$, not in $\nu$. If (as occasionally happens) both deviations are in the same direction then special care is needed, as described in the section on ‘flipped’ distributions, Section 2.1.1.
@@ 1.2 Beyond the Gaussian
for a measurement $x$ of some value $a$. This is what is encoded in any reported result of the form $\hat{a}=x\pm\sigma$. When a result is reported as $\hat{a}={x}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ then this is making a clear analogy with the Gaussian form. But this similarity conceals two important questions which cannot be evaded.
@@
The difference between rms spread ‘pdf errors’ and confidence interval ‘likelihood errors’ requires clarification. They both refer to uncertainties, the former from random probability and the latter from statistical inference. The Poisson process provides a simple illustration (though for a discrete probability distribution rather than a continuous pdf): the former describes the uncertainty of predicting results from a particular true mean, the latter the uncertainty of inferring the true value from a particular result.
Pdf errors arise especially through measurements and their errors in experimental work. At high school and university the physicist is firmly taught never to quote an experimental result as a single number $x$, but always to give its error $\sigma$. That $\sigma$ is the standard deviation of the Gaussian distribution, as given in (1), from which this value is a sample. If, for better accuracy, the mean of $N$ measurements is quoted then it is accompanied by the standard error $\sigma/\sqrt{N}$ which is the standard deviation of the pdf of $\overline{x}$. The well-known formula for the combination of errors, given below as Equation 2, known to statisticans as the delta method, builds on this. It describes the combination of measurement errors: if $I$ is sampled from a Gaussian with standard deviation $\sigma_{I}$, and $V$ with $\sigma_{V}$, then the resistance measurement $R=V/I$ is sampled from a Gaussian for which (assuming the function can legitimately be linearised) $\sigma_{R}^{2}=\sigma_{V}^{2}/I^{2}+V^{2}\sigma_{I}^{2}/I^{4}$. The quoted result, $V/I\pm\sigma_{R}$, is saying the value has been drawn from a Gaussian distribution of unknown mean and standard deviation $\sigma_{R}$.
Such measurements can be converted to confidence intervals using the Neyman confidence belt, as described in Section 40 of the Particle Data Group handbook [12]. For Gaussians this is trivial. In frequentist usage, to say “I have measured $a$ to be $x\pm 3$.” means “I have a measurement $x$ from a device which returns values distributed about the true value $a$ according to a Gaussian distribution with a standard deviation of 3: I therefore assert that $x$ falls within the interval $[a-3,a+3]$ with probability 68%, which in the Gaussian case can be turned into saying that, given $x$, the interval $[x-3,x+3]$ covers the true value $a$ with 68% confidence. (It does not mean “The true value of $a$ lies within within $x-3$ and $x+3$ with 68% probability.”, that would be a Bayesian credible interval, though this subtle distinction may not always be appreciated.)
If the pdf is not Gaussian – and for asymmetric errors this must be the case – then this trivial equivalence breaks. ‘Pdf errors’ describe the standard deviation and ‘likelihood errors’ the 68% central confidence interval. To take an artificially exaggerated but hopefully illuminating example, suppose a measuring device has a 50% chance of recording a measurement below the true value and does so according to a half-Gaussian with $\sigma=1$, and a 50% chance of recording a measurement above the true value, according to a half-Gaussian with $\sigma=5$. (This is the dimidiated Gaussian, described later.) If it measures $x=11.7$ this would be written as ${11.7}^{+5.0}_{-1.0}$, to describe the pdf. However the Neyman construction for the confidence belt would be ${11.7}^{+1.0}_{-5.0}$: if 11.7 is a downward fluctuation it cannot have come very far from the true value, whereas an upward fluctuation may have been produced by a true value ome distance away. This is shown in Figure 2.
Figure 2: The Neyman confidence belt referred to in the text, showing how a positive pdf asymmetry ${11.7}^{+5.0}_{-1.0}$ matches a negative likelihood asymmetry ${11.7}^{+1.0}_{-5.0}$. The shaded areas display 68% confidence regions of the pdf. This is an artificial example to make the point clear: real life is more complicated.
It is tempting to suggest that any result presented with an asymmetric error should denote a confidence-level, and there is a case for doing so for final results. However in an analysis there will be intermediate values used to construct a final result using addition-in-quadrature (with the aid of the combination-of-errors formula), and this is done in the framework of pdfs and their moments. Variances add under convolution, but 68% central regions do not, requiring more complicated meta-analysis procedures. In a typical particle physics analysis, there are many intermediate steps between the raw data and the final quoted figure, and typically for these the errors quoted are properties of the pdf.
@@
If we are considering a pdf $p(x;a)$ then we can integrate it to obtain the spread (as given by the variance) of the measured $x$ for a given value of $a$ and the ‘errors’ are expression of that spread. It tells us nothing about $a$, as that is given. If we are dealing with a likelihood function $L(a;x)$ we can get the ‘best’ estimate $\hat{a}$ and take the 68% CL central region from the $\Delta\ln L=-{1\over 2}$ points, [^1] This is only an approximation to the Neyman construction, but this point is generally not considered. For a discussion see Reference [13]. but we can say nothing more about $x$. It is true that a pdf can give 68% central limits, but in the language of the Neyman confidence belt they are the horizontal lines constructed on the plot it shows in Figure 2, not the vertical ones from which the confidence interval for a result is read off.So an ‘asymmetric error’ may refer to the variance and skewness of a pdf, or to the confidence region for a likelihood. Both cases arise in practice. We will therefore consider asymmetric pdfs and the variances (and higher moments) of $x$, and also, separately, asymmetric likelihoods and the confidence regions for $a$.
@@
     Usage varies in the literature, but within this paper we use the variable names $x$ and sometimes $y$ to refer to results of measurements, and $a$ to refer to an ideal ‘true’ value. Thus pdfs are considered as functions of $x$ and likelihoods as functions of $a$. Alternative names may be used in particular examples, as when $R$ and $\nu$ (for ‘Result’ and ‘nuisance’) are used in the OPAT plot of Figure 1 rather than $\hat{a}$ and $x$.
@@ 1.2.1 Combination of Errors
This simple formula is where most physicists first meet statistics, in the context of experimental errors and uncertainties. For some function $u=u(x,y)$ where $x$ and $y$ are statistically independentwith errors that are small (so that a linear approximation is valid),
@@
We need to extend it to cope with $\sigma^{+}_{x},\sigma^{-}_{x},\sigma^{+}_{y},\sigma^{-}_{y}$ and thus $\sigma^{+}_{u},\sigma^{-}_{u}$.
In introductory texts this formula is used in problems like determining the speed from the distance travelled in a given time using $v=x/t$, or the acceleration due to gravity from the length and period of a simple pendulum using $g=4\pi^{2}L/T^{2}$. In more advanced experimental work it is used in instances like calculating a branching ratio, $Br={N-b\over\eta N_{T}}$, from some number $N$ of observed decays into the channel of interest out of a total of $N_{T}$ decays,with a background $b$ and an efficiency $\eta$, or in fitting a straight line (or some more sophisticated function) to a set of values $\{x_{i},y_{i}\pm\sigma_{i}\}$.
@@
where the two random variables $X_{1}$ and $X_{2}$ are supposed independent. If the pdfs are Gaussian, as shown in the left hand plot of Figure 3, where the dotted lines are the lines of constant $x_{2}-x_{1}$, then the integral gives a Gaussian for $\ell$. This Gaussian has standard deviation $\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}$, as variances add. This is what is expressed by Equation (2).
Figure 3: Combination of errors: $\ell=(2.5\pm 1.0)-(-2.5\pm 1.0)$ using pdfs (left) and likelihoods (right).
@@ 1.2.2 Combination of Results
But it is also possible that one has a set of results presented as pdfs: the $\hat{a}_{i}$ are estimators of the true $a$, they are functions of the data $x_{i}$ and are therefore described by pdfs. If they are unbiassed then any weighted sum $\sum_{i}w_{i}\hat{a}_{i}$ is an unbiased estimate, and one has the freedom to choose the $w_{i}$ to minimise the variance on this, using the combination of errors formula, Equation (2). For Gaussian errors this gives the well known prescription $w_{i}={1/\sigma_{i}^{2}\over\sum_{j}1/\sigma_{j}^{2}}$: the weight is proportional to the inverse square of the error. Hence for Gaussian errors the combination results using pdfs and using likelihoods gives the same answer, even though the meaning of the answer is subtly different. Again, for non-Gaussian errors this is not the case and we must consider both separately.
1.3 Summary on combination-of-error and combination-of-results
The distinction between combination of errors and combination of results turns out to require clarification, perhaps because both are considered as ‘the error analysis’. The fundamental difference is that, in the first case, we want to quantify the precision of a combination, which involves different quantities, while in the second case, we compute the precision of an updated measurement of the same quantity, we obtain by combining different measurements of the same quantity. In the first case, precision goes down, because uncertainty adds. In the second case, precision goes up because we measure something with more data. (In statistics, the same name is kept for the precision, but different names are used for the outcome: measurement in the first case, estimate in the second.)
Combination of errors fits naturally into the framework of pdf errors, though combination of results is perfectly possible where the result is a weighted sum of measured values. Conversely combination-of-results emerges very naturally from likelihoods, although combination of errors is perfectly possible through profiling.
For clarification the distinctions are shown in Table 1.
Combination of/Through PDF (Section 2) LIKELIHOOD (Section 3) Notes
ERROR
(Sections 2.2
and 3.4)
Want to quantify variability of a measurement obtained by combining measurements of different quantities.
$X\sim p_{x}(x;a_{x},\sigma_{x}^{+},\sigma_{x}^{-})$
$Y\sim p_{y}(y;a_{y},\sigma_{y}^{+},\sigma_{y}^{-})$
Parameter of interest: $\psi=a_{x}/a_{y}$, fitted by $\hat{\psi}=x/y$ (for example).
Uncertainty of $\hat{\psi}$ is quantified through
$U=X/Y\sim p_{u}(u;\psi,\sigma_{u}^{+},\sigma_{u}^{-}).$
Bayesian view:
$A_{x}\sim p_{1}(a_{x}\mid x,\sigma_{x}^{+},\sigma_{x}^{-})$
$A_{y}\sim p_{2}(a_{y}\mid y,\sigma_{y}^{+},\sigma_{y}^{-})$
combined to yield
$U=A_{x}/A_{y}\sim p_{u}(u\mid x,y,\sigma_{u}^{+},\sigma_{u}^{-})$
$L(a_{1};x)\quad\text{and}\quad L(a_{2};y)$
$\Downarrow\newline L(a_{1},a_{2};x,y)=L(\psi,\lambda;x,y)$,
with $\psi=a_{1}/a_{2}$ parameter of interest (for example) and $\lambda$ any function of $(a_{1},a_{2})$ such that $(\psi,\lambda)$ is a reparametrization of $(a_{1},a_{2})$
$\bullet$ The parameter of interest $\psi$ is a combination of other model parameters (say, $a_{1}$ and $a_{2}$).
$\bullet$ $n$ represents the number of random variables (here, $X$ and $Y$) which provide information (here on $a_{1}$ and $a_{2}$). These are then combined to provide information on $\psi$.
$\bullet$ Uncertainty goes up!
$\bullet$ The procedure is usually based on PDFs (where $\sigma$ represents an rms error).
$\bullet$ If based on LIKELIHOODS ($\sigma$ then typically represents the spread of an 68% CI), this requires profiling out the nuisance parameters $\lambda$ which complete the reparametrization from $(a_{1},a_{2})$ to $(\psi,\lambda)$.
RESULT
(Sections 2.4
and 3.2)
Want to quantify confidence of measurement obtained by combining measurements of the same quantity.
Weighted average (meta-analysis) $\hat{a}=\sum_{i}w_{i}\hat{a}_{i}$, with $\sum w_{i}=1$ Combine
${\hat{a}(x_{1})}^{+\sigma^{+}(x_{1})}_{-\sigma^{-}(x_{1})}$  from $L(a;x_{1})$
${\hat{a}(x_{2})}^{+\sigma^{+}(x_{2})}_{-\sigma^{-}(x_{2})}$  from $L(a;x_{2})$
to yield
${\hat{a}(x)}^{+\sigma^{+}(x)}_{-\sigma^{-}(x)}$  from $L(a;x),$
with $x=(x_{1},x_{2})$.
$\bullet$ There is an unique parameter of interest $a$.
$\bullet$ $n$ is the number of observations available on $a$.
$\bullet$ Uncertainty goes down!
$\bullet$ The procedure is usually based on LIKELIHOODS (where $\sigma$ represents the spread of an 68% CI).
$\bullet$ It can be based on PDFs (where $\sigma$ represents an rms error) using meta-analysis.
$\bullet$ Check for goodness-of-fit!
Notes $\bullet$ $\sigma$ measures variability (e.g. precision of measurement device).
$\bullet$ Approximation needs be normalized (to yield a PDF).
$\bullet$ $\sigma$ measures the confidence region (e.g. as expressed by Neyman confidence belt).
$\bullet$ Approximation needs be “nearly” parabolic (to work well).
Table 1: Combination of errors and results using pdf errors and likelihood errors.
@@ 2.1 Modelling pdf errors
    the quantiles. The median $M$ is the middle quantile $x_{0.50}$, and the 68% (one sigma) central confidence probability region $[x_{0.16},x_{0.84}]$ can be written as $[M-{\sigma^{-}},M+{\sigma^{+}}]$. This can be expressed in the form ${M}^{+{\sigma^{+}}}_{-{\sigma^{-}}}$ and will be referred to as “the quantile parameters” in what follows. (The region is actually bounded by the values of the cumulative Gaussian $\Phi(-1)$ and $\Phi(+1)$: though here and elsewhere only 2 significant figures are shown the exact one sigma region is always meant.)
@@
It might be thought that the mean could be used as an alternative to the median in combination with the bounds of the 68% central confidence region, writing it as $[x_{0.16},x_{0.84}]=[\mu-{\sigma^{-}},\mu+{\sigma^{+}}]$. However, this parameterisation has an unhelpful behaviour in some models, as will be noted in Sections A.8 and A.9, and it will not be pursued further.
@@
Figure 4: The dimidiated and distorted Gaussians for ${5.00}^{+1.0}_{-0.9}$ and ${5.00}^{+0.85}_{-1.15}$.